# FPT+: A Parameter and Memory Efficient Transfer Learning Method for High-resolution Medical Image Classification

### v1

@@ 4.2.1 Experiment Setup
For compared methods, in the context of LoRA (Hu et al., 2022) and AdaLoRA (Zhang et al., 2023a), we set the scale factor $\alpha$ to 16 and add LoRA layers in the weights of query and value layers in ViTs. We reproduce prompt tuning following the design of VPT-deep (Jia et al., 2022), where the number of prompts is set to 16. To ensure fair comparisons across methods, grid searching is utilized to determine the optimal set of hyperparameters, including the learning rate and the weight decay, based on performance on the validation set. All methods and the side network of FPT+ are trained using the same data augmentations, including cropping, horizontal flipping, and color distortion.
4.2.2 Evaluation Metrics
We use the Area Under the Receiver Operating Characteristic Curve (AUC) to evaluate the classification performance on each dataset. To evaluate the performance and efficiency balance, we adopt the metrics proposed in our previous work (Huang et al., 2024a), which extend from the performance-efficiency metric (Li et al., 2022; He et al., 2023b). Specifically, two metrics, namely performance-parameter-efficiency (PPE) and performance-memory-efficiency (PME), are utilized to assess the impact of the number of learnable parameters and GPU memory consumption on efficiency during transfer learning, respectively. They are defined as
@@ 4.2.2 Evaluation Metrics
Table 2: Comparison results with state-of-the-art PETL methods using a ViT-L backbone with supervised pre-trained weights, evaluated at different input resolutions. The ’/’ symbol indicates out-of-memory conditions on a 48GB GPU. The average AUC is computed across all eight classification tasks.
Memory Usage Average AUC
@@
Full FT 7816 12168 26470 / / 83.94 86.96 87.23 / /
Linear probing 1640 1884 2778 5066 9716 77.46 82.99 84.38 84.62 84.77
Prompt-tuning (Jia et al., 2022) 3128 7602 20392 / / 79.82 83.90 84.95 / /
Attention-tuning (Touvron et al., 2022) 4340 8710 21516 / / 80.93 82.21 83.89 / /
Adapter (Chen et al., 2022) 3248 7608 20394 / / 79.30 81.70 82.88 / /
BitFit (Ben-Zaken et al., 2022) 3058 7186 20074 / / 82.85 85.46 86.38 / /
LoRA (Hu et al., 2022) 3152 7810 21190 / / 80.13 82.96 83.95 / /
AdaLoRA (Zhang et al., 2023a) 3142 7806 21056 / / 82.17 85.52 86.46 / /
FPT (Huang et al., 2024a) 2546 2724 3152 3730 4456 80.72 84.34 85.75 86.85 87.28
FPT+ (Ours) 570 630 718 862 1050 81.74 85.60 86.88 87.87 88.14
4.3 Comparisons with State-of-the-art
@@
We compare FPT+ against full fine-tuning, linear probing, and state-of-the-art (SOTA) PETL approaches, including our previous work FPT. In full fine-tuning, all parameters are learnable when training on downstream tasks. Although full fine-tuning often represents the performance’s upper bound for fine-tuning, the relative scarcity of training samples in medical image datasets may lead to overfitting and consequent performance degradation. Linear probing emerges as an efficient strategy to adapt pre-trained models to downstream tasks by solely training a new task-specific head on top of the pre-trained model. Additionally, we compare FPT+ with popular SOTA PETL methods applicable to vision tasks.
@@
The performance comparisons are tabulated in Table 1. Full fine-tuning, while providing competitive performance across all datasets, incurs the highest computational cost, with a 100% parameter ratio and substantial memory usage (23,128 MB). This makes it less feasible for resource-constrained settings despite its strong performance. Linear probing, on the other hand, shows low parameter usage (0.01%) and memory consumption (3,416 MB). However, this method generally underperforms. Other PETL methods are more balanced between parameter efficiency and performance. For example, LoRA and AdaLoRA, with 0.68% and 0.52% parameter ratios respectively, show competitive performance across various datasets.
@@
Although these methods achieve reductions in the number of learnable parameters, their memory consumption remains high. For instance, adapter has the lowest memory consumption but still requires 19,360 MB to fine-tune the LPM. In contrast, our proposed method, FPT+, outperforms other PETL methods in both performance and efficiency, achieving the best performance among PETL methods for both pre-trained weights and the best performance-efficiency balance (PPE and PME) among all methods. Specifically, FPT+ exhibits remarkable parameter efficiency with only 1.03% of learnable parameters and significantly lower memory consumption (736 MB, 3.18% compared to full fine-tuning), without compromising performance. Our previous work, FPT, already shows significant improvements in parameter efficiency and performance. On top of FPT, FPT+ further enhances these metrics, achieving higher performance and reduced memory consumption (736 MB vs. 1,824 MB).
@@
The comparison results emphasize the effectiveness of FPT+ in balancing performance, parameter efficiency, and memory consumption. Its design ensures efficient high-resolution medical image classification, making it a promising solution for advancing medical diagnostics and related fields. The generalizability and robustness of FPT+ are further highlighted by its effective use of both supervised and DINO pre-trained weights, making it adaptable to different training paradigms and practical for a wide range of medical image classification tasks.
4.4 Evaluation with A Larger Pre-trained Model
@@
To further investigate the generalizability of FPT+ and highlight its significant advantage in memory saving, we evaluate its performance and efficiency by comparing it with SOTA methods using a ViT-L backbone with supervised pre-trained weights across different input resolutions. The superiority of FPT+ becomes even more evident when using a larger model. As shown in Table 2, most PETL methods encounter out-of-memory issues on a 48GB GPU when the input resolution exceeds 384 $\times$ 384, limiting their capability to improve performance through higher resolutions. In contrast, FPT, FPT+, and linear probing, with better memory efficiency, can run at higher resolutions and exhibit greater performance potential.
@@
Full fine-tuning achieves its best performance with an average AUC of 87.23 at its highest feasible resolution of 384 $\times$ 384. Similarly, the PETL method AdaLoRA achieves its best performance with an average AUC of 86.48, also at the resolution of 384 $\times$ 384. These methods are constrained by memory limitations, which largely restrict their feasibility to operate at higher resolutions. Although linear probing can run at a resolution of 640 $\times$ 640, its performance remains relatively low, with an average AUC of 84.77 at the resolution of 640 $\times$ 640. In contrast, FPT+ not only achieves the best performance among all PETL methods at the resolution of 384 $\times$ 384 with an average AUC of 86.88 but also runs efficiently at the resolution of 640 $\times$ 640, achieving the highest overall performance with an average AUC of 88.14. These results demonstrate the superior scalability of FPT+ when handling higher resolutions.
@@
One of the key advantages of FPT+ is its exceptional memory efficiency. At a resolution of 640 $\times$ 640, FPT+ has a memory usage of 1,050 MB, which is even lower than the memory usage of other methods at a resolution of 128 $\times$ 128. Such memory efficiency enhances the feasibility of deploying FPT+ and enables the exploration of higher input resolutions to potentially achieve even better performance.
@@
4.5 Contribution of Each Layer
@@
Since higher layers of a ViT-based LPM are more likely to contain high-level information for image understanding (Raghu et al., 2021), we conjecture that not all layers of the side network are necessary for high performance. To investigate the contribution of each layer to downstream tasks, we visualize the average attention score that other tokens pay to the fine-grained prompts at each layer of FPT. As shown in Fig. 5, the fusion modules pay more attention to the fine-grained prompts in higher layers; the average attention score of layers 7-12 is 0.315, while that of layers 1-6 is 0.180. This suggests that leveraging information from the higher layers is more critical for downstream tasks, allowing for a more efficient design of the side network by focusing on the key layers.
@@
4.6 Ablation Study
In this section, we systematically investigate the impact of each component in FPT+. Unless otherwise specified, the presented performance metric is the average AUC across eight datasets using the pre-trained weights in a supervised manner, and the configuration of FPT+ remains the same as that described in Section 4.2.1.
4.6.1 Impact of Components
@@
4.6.2 Input Resolutions
To reduce memory consumption, we propose asymmetric inputs for the LPM and the side network, where the input for the side network is a down-scaled version of that for the LPM. In this section, we investigate the impact of the input resolution for the side network. We train FPT+ using different input resolutions for the side network while maintaining the input resolution for the LPM at $512\times 512$. As shown in Table 3, the results indicate that a resolution of $128\times 128$ yields optimal performance while maintaining memory efficiency. We conjecture that the poor performance observed with a higher down-scaled resolution could potentially be attributed to the lengthy input sequence in the side network, which distracts attention from the fine-grained prompts. Notably, even using an input resolution of $64\times 64$ for the side network achieves better performance compared to other PETL methods, demonstrating that the proposed fine-grained prompts and fusion modules can effectively convey fine-grained information for diagnostics in the side network.
4.6.3 Number of Fine-grained Prompts
The number of fine-grained prompts can be significantly fewer than the total tokens from the input images presented in both the pre-trained model and the side network, thereby reducing the memory consumption. We here explore the influence of the number of fine-grained prompts. As depicted in Table 4, with the increase in the number of prompts, the performance improves and the best performance is achieved at 16 prompts, which number is sufficient to summarize fine-grained information and convey it to the side network. We also notice that even with only 4 prompts, the performance is comparable to other PETL methods, demonstrating the effectiveness of FPT+’s fusion modules in obtaining pre-trained knowledge from the LPM.
4.6.4 Ratio of Important Tokens
Table 3: FPT+’s performance and efficiency with different input resolutions for the side network.
Input Resolution Avg. AUC Mem. PME
@@ 4.6.4 Ratio of Important Tokens
256 86.32 1342 84.23
512 84.78 8704 73.80
Table 4: FPT+’s performance and efficiency with different numbers of fine-grained prompts.
# Prompt Avg. AUC Param. Mem. PPE PME
@@
128 86.49 1.61 1388 85.89 84.33
@@ 5 Discussion
High-resolution image classification is not the only field that benefits from this memory efficiency. FPT+ can also be applied to ViT variants with smaller patch sizes. ViTs with smaller patch sizes generally have more powerful abilities to extract fine-grained semantics. However, smaller patch sizes result in longer input sequences for the same input resolution. FPT+ may be a good solution to attain the best balance. The memory efficiency of FPT+ also allows for larger batch sizes during training, which can accelerate the training process and potentially improve the stability and performance of transfer learning. Additionally, there are settings in natural image analysis that also require high resolution, such as satellite imagery (Adegun et al., 2023). The generalizability of FPT+ to natural images remains unexplored, which is one of our future works.
6 Conclusion
In this paper, we extend our previous work Fine-grained Prompt Tuning (FPT) (Huang et al., 2024a) to Fine-grained Prompt Tuning plus (FPT+). FPT+ significantly improves the memory efficiency over existing PETL methods, particularly in the high-resolution context commonly encountered in medical image analysis. To reduce memory consumption, we first enhance the design of side tuning with asymmetric input. To effectively adapt pre-trained knowledge from the LPM, we introduce fine-grained prompts and fusion modules to summarize and process intermediate features of the LPM and effectively convey them to the side network. To further reduce memory utilization, we propose the selection of important tokens and the preloading of pre-trained features. By integrating all these components, we present a parameter and memory-efficient transfer learning method for high-resolution medical image classification. The effectiveness of FPT+ is successfully established across eight medical image datasets of diverse sizes, modalities, and complexities. Experimental results show that FPT+ outperforms all other SOTA PETL methods while maintaining the best parameter and memory efficiency.
CRediT authorship contribution statement
Yijin Huang: Writing – original draft, Conceptualization, Investigation, Methodology, Validation, Visualization. Pujin Cheng: Investigation, Validation, Writing – review and editing. Roger Tam: Resources, Supervision, Writing – review and editing. Xiaoying Tang: Funding acquisition, Resources, Supervision, Writing – review and editing.
Declaration of competing interest
The authors have no competing interests to declare that are relevant to the content of this article.
Acknowledgements
This study was supported by the National Key Research and Development Program of China (2023YFC2415400); the National Natural Science Foundation of China (62071210); the Guangdong Basic and Applied Basic Research (2024B1515020088); the Shenzhen Science and Technology Program (RCYX20210609103056042); the Shenzhen Science and Technology Innovation Committee (KCXFZ2020122117340001); the Guangdong Basic and Applied Basic Research (2021A1515220131); the High Level of Special Funds (G030230001, G03034K003).

### v2

@@ IV-B1 Experiment Setup
For compared methods, LoRA [22] and AdaLoRA [51] use a scale factor ($\alpha$) of 16, with LoRA layers applied to the query and value weights in the ViT. Prompt tuning is reproduced following the VPT-deep design [27], with 16 prompts. To ensure fair comparisons across methods, grid searching is utilized to determine the optimal set of hyperparameters, including the learning rate and the weight decay, based on performance on the validation set. All methods and the side network of FPT+ are trained using the same data augmentations, including cropping, horizontal flipping, and color distortion.
IV-B2 Evaluation Metrics
We use the Area Under the Receiver Operating Characteristic Curve (AUC) to evaluate the classification performance on each dataset. To evaluate the performance and efficiency balance, we adopt the metrics proposed in our previous work [23], which extend from the performance-efficiency metric [32, 20]. Specifically, two metrics, namely performance-parameter-efficiency (PPE) and performance-memory-efficiency (PME), are utilized to assess the impact of the number of learnable parameters and GPU memory consumption on efficiency during transfer learning, respectively. They are defined as
@@ IV-B2 Evaluation Metrics
TABLE II: Training efficiency comparisons across different PETL methods on the Messidor2 dataset.
FLOPs (GB) $\downarrow$
Total Training
time (s)
$\downarrow$
Throughput
(img/s)
$\uparrow$
Peak Power
Usage (W)
$\downarrow$
Avg. Power
Usage (W)
$\downarrow$
Full FT 87.84 1135.91 18.48 295.62 280.92
Linear 87.84 416.72 51.06 286.61 278.59
Prompt-tuning [27] 88.30 946.55 22.19 296.69 286.52
Attention-tuning [45] 87.84 1005.62 20.89 291.76 281.79
Adapter [7] 89.68 954.51 21.93 297.30 288.30
BitFit [2] 87.84 966.69 21.75 291.40 279.66
LoRA [22] 107.89 1014.68 20.62 295.85 288.47
AdaLoRA [51] 107.28 1027.17 20.44 292.00 280.56
Preloading 87.84 31.76 32.81 221.812 217.74
FPT 0.28 154.23 171.83 166.54 183.10
FPT+ 0.06 90.61 357.37 114.38 107.60
TABLE III: Comparison results with state-of-the-art PETL methods using a ViT-L backbone with supervised pre-trained weights, evaluated at different input resolutions. The ’/’ symbol indicates out-of-memory conditions on a 48GB GPU. The average AUC is computed across all eight classification tasks.
Memory Usage $\downarrow$ Average AUC $\uparrow$
@@
Full FT 7,816 12,168 26,470 / / 83.94 86.96 87.23 / /
Linear probing 1,640 1,884 2,778 5,066 9,716 77.46 82.99 84.38 84.62 84.77
Prompt-tuning [27] 3,128 7,602 20,392 / / 79.82 83.90 84.95 / /
Attention-tuning [45] 4,340 8,710 21,516 / / 80.93 82.21 83.89 / /
Adapter [7] 3,248 7,608 20,394 / / 79.30 81.70 82.88 / /
BitFit [2] 3,058 7,186 20,074 / / 82.85 85.46 86.38 / /
LoRA [22] 3,152 7,810 21,190 / / 80.13 82.96 83.95 / /
AdaLoRA [51] 3,142 7,806 21,056 / / 82.17 85.52 86.46 / /
FPT [23] 2,546 2,724 3,152 3,730 4,456 80.72 84.34 85.75 86.85 87.28
FPT+ (Ours) 570 630 718 862 1,050 81.74 85.60 86.88 87.87 88.14
IV-C Comparisons with State-of-the-art
@@
We compare FPT+ against full fine-tuning, linear probing, and SOTA PETL approaches, including our previous work FPT. In full fine-tuning, all parameters are learnable when training on downstream tasks. Although full fine-tuning often represents the performance’s upper bound for fine-tuning, the relative scarcity of training samples in medical image datasets may lead to overfitting and consequent performance degradation. Linear probing emerges as an efficient strategy to adapt pre-trained models to downstream tasks by solely training a new task-specific head on top of the pre-trained model. Additionally, we compare FPT+ with popular SOTA PETL methods applicable to vision tasks.
@@
The performance comparisons are tabulated in Table I. Full fine-tuning presents strong performance across all datasets but comes with the highest computational cost, requiring a 100% parameter ratio and significant memory usage (23,128 MB), making it impractical for resource-constrained settings. Linear probing, on the other hand, shows low parameter usage (0.01%) and memory consumption (3,416 MB). However, this method generally underperforms. Other PETL methods are more balanced between parameter efficiency and performance. For example, LoRA and AdaLoRA, with 0.68% and 0.52% parameter ratios respectively, show competitive performance across various datasets.
@@
Although these aforementioned methods achieve reductions in the number of learnable parameters, their memory consumption remains high. For instance, Adapter has the lowest memory consumption but still requires 19,360 MB to fine-tune the LPM. In contrast, our proposed method, FPT+, outperforms other PETL methods in both performance and efficiency, achieving the best performance among PETL methods for both pre-trained weights and the best performance-efficiency balance (PPE and PME) among all methods. Specifically, FPT+ exhibits remarkable parameter efficiency with only 1.03% of learnable parameters and significantly lower memory consumption (736 MB, 3.18% compared to full fine-tuning), without compromising performance. Our previous work, FPT, already shows significant improvements in parameter efficiency and performance. On top of FPT, FPT+ further enhances these metrics, achieving higher performance and reduced memory consumption (736 MB vs. 1,824 MB).
@@
The comparison results emphasize the effectiveness of FPT+ in balancing performance, parameter efficiency, and memory consumption. Its design ensures efficient high-resolution medical image classification, making it a promising solution for advancing medical diagnostics. The generalizability and robustness of FPT+ are further highlighted by its effectiveness with both supervised and DINO pre-trained weights, making it adaptable to different pre-training paradigms and practical for a wide range of medical image classification tasks.
IV-D Training Efficiency
To comprehensively evaluate the efficiency of PETL methods, we further assess several metrics related to training cost, including computational cost (FLOPs), total training time, training time per epoch, throughput, and power usage (both peak and average). These training efficiency metrics are evaluated using the Messidor2 dataset, containing 1,748 samples with a resolution of 512 × 512, and employing ViT-B as the backbone. Table II provides a detailed comparison of these metrics across various methods.
As shown in Table II, FPT+ demonstrates remarkable training efficiency with computational cost of only 0.06 GB FLOPs, being significantly lower than full fine-tuning (87.84 GB) and PETL methods like LoRA (107.89 GB). FPT+ also excels in terms of throughput, processing 357.37 images per second, which is a significant improvement over other PETL methods like LoRA (20.62 images/second) and Prompt-tuning (22.19 images/second). Compared to FPT, FPT+ doubles the throughput, showing its enhanced efficiency. With these reductions in the computational cost, FPT+ achieves a total training time of 90.61 seconds, which is more than 12 times faster than full fine-tuning (1135.91 seconds). FPT+ is also energy-efficient, achieving an average power usage of 107.60 W and a peak power usage of 114.38 W, staying far below the average of 280 W and the peak usage of over 290 W for other PETL methods on GPUs with a 300 W power limit.
Additionally, FPT+ incorporates a preloading step, which extracts and pre-stores features from the frozen LPM before training. This step, taking only 31.76 seconds, adds negligible overhead compared to the substantial training costs of other fine-tuning methods.
These results establish FPT+ as a highly efficient PETL method in terms of computational, temporal, and energy resources, making it a practical and effective solution for high-resolution medical image classification tasks, especially in resource-constrained settings.
IV-E Evaluation with A Larger Pre-trained Model
@@
To further investigate the generalizability of FPT+ and highlight its significant advantage in memory saving, we evaluate its performance and efficiency by comparing it with SOTA methods using a ViT-L backbone with supervised pre-trained weights across different input resolutions. The superiority of FPT+ becomes even more evident when using a larger model. As shown in Table III, most PETL methods encounter out-of-memory issues on a 48GB GPU when the input resolution exceeds 384 $\times$ 384, limiting their capability to improve performance through higher resolutions. In contrast, FPT, FPT+, and linear probing, with better memory efficiency, can run at higher resolutions and exhibit greater performance potential.
@@
Full fine-tuning achieves its best performance with an average AUC of 87.23, at its highest feasible resolution of 384 $\times$ 384. Similarly, the PETL method AdaLoRA achieves its best performance with an average AUC of 86.48, also at the resolution of 384 $\times$ 384. These methods are constrained by memory limitations, which largely restrict their feasibility to operate at higher resolutions. Although linear probing can run at a resolution of 640 $\times$ 640, its performance remains relatively low, with an average AUC of 84.77 at the resolution of 640 $\times$ 640. In contrast, FPT+ not only achieves the best performance among all PETL methods at the resolution of 384 $\times$ 384 with an average AUC of 86.88 but also runs efficiently at the higher resolution of 640 $\times$ 640, achieving the highest overall performance with an average AUC of 88.14. These results demonstrate the superior scalability of FPT+ when handling higher resolutions.
@@
One of the key advantages of FPT+ is its exceptional memory efficiency. At a resolution of 640 $\times$ 640, FPT+ has a memory usage of 1,050 MB, which is even lower than the memory usage of other methods at a resolution of 128 $\times$ 128. Such memory efficiency enhances the feasibility of deploying FPT+ and enables the exploration of higher input resolutions to potentially attain even better performance.
@@
IV-F Contribution of Each Layer
@@
Since higher layers of a ViT-based LPM are more likely to contain high-level information for image understanding [39], we conjecture that not all layers of the side network are necessary for high performance. To investigate the contribution of each layer to downstream tasks, we visualize the average attention score that other tokens pay to the fine-grained prompts at each layer of FPT. As shown in Fig. 6, the fusion modules pay more attention to the fine-grained prompts in higher layers; the average attention score of layers 7-12 is 0.315, while that of layers 1-6 is 0.180. This suggests that leveraging information from higher layers is more critical for downstream tasks, allowing for a more efficient design of the side network by focusing on key layers.
@@
IV-G Ablation Study
In this section, we systematically investigate the impact of each component in FPT+. Unless otherwise specified, the presented performance metric is the average AUC across eight datasets using the pre-trained weights in a supervised manner, and the configuration of FPT+ remains the same as that described in Section IV-B1.
IV-G1 Impact of Components
@@
IV-G2 Input Resolutions
To reduce memory consumption, we propose asymmetric inputs for the LPM and the side network, where the input for the side network is a down-scaled version of that for the LPM. In this section, we investigate the impact of the input resolution for the side network. We train FPT+ using different input resolutions for the side network while maintaining the input resolution for the LPM at $512\times 512$. As shown in Table IV, the results indicate that a resolution of $128\times 128$ yields optimal performance while maintaining memory efficiency. We conjecture that the poor performance observed with a higher down-scaled resolution could potentially be attributed to the lengthy input sequence in the side network, which distracts attention from the fine-grained prompts. Notably, even using an input resolution of $64\times 64$ for the side network achieves better performance compared to other PETL methods, demonstrating that the proposed fine-grained prompts and fusion modules can effectively convey fine-grained information for diagnostics in the side network.
TABLE IV: FPT+’s performance and efficiency with different input resolutions for the side network.
Input Resolution Avg. AUC $\uparrow$ Mem. $\downarrow$ PME $\uparrow$
@@ IV-G2 Input Resolutions
256 86.32 1,342 84.23
512 84.78 8,704 73.80
IV-G3 Number of Fine-grained Prompts
The number of fine-grained prompts can be significantly fewer than the total tokens from the input images presented in both the pre-trained model and the side network, thereby reducing the memory consumption. We here explore the influence of the number of fine-grained prompts. As depicted in Table V, with the increase in the number of prompts, the performance improves and the best performance is achieved at 16 prompts, which number is sufficient to summarize fine-grained information and convey it to the side network. We also notice that even with only 4 prompts, the performance is comparable to other PETL methods, demonstrating the effectiveness of FPT+’s fusion modules in obtaining pre-trained knowledge from the LPM.
TABLE V: FPT+’s performance and efficiency with different numbers of prompts for the side network.
# Prompt Avg. AUC $\uparrow$ Param. $\downarrow$ Mem. $\downarrow$ PPE $\uparrow$ PME $\uparrow$
@@ IV-G3 Number of Fine-grained Prompts
128 86.49 1.61 1,388 85.89 84.33
IV-G4 Reduction Factor
The reduction factor $k$ in FPT+ determines the hidden dimension of the side network. As shown in Table VI, we assess the impact of different reduction factors on FPT+’s performance and efficiency. A reduction factor of 8 achieves the highest average AUC (87.12) with minimal memory usage (736 MB) and only 1.03% of learnable parameters. Increasing the reduction factor to 16 slightly reduces memory usage but results in a drop in AUC (85.92), while lower factors (e.g., 2 and 4) increase resource consumption without significant gains in performance. We observe that a larger hidden dimension in the side network (lower reduction factor) does not improve performance as it increases the training complexity of the randomly initialized side network. A reduction factor of 8 achieves the optimal balance for high-resolution medical image classification tasks.
TABLE VI: FPT+’s performance and efficiency with different reduction factors.
Reduction Factor Avg. AUC $\uparrow$ Param. $\downarrow$ Mem. $\downarrow$ PPE $\downarrow$ PME $\downarrow$
2 86.32 11.55 1,170 82.32 84.49
4 86.74 3.40 894 85.49 85.32
8 87.12 1.03 736 86.73 85.94
16 85.92 0.37 724 85.78 84.78
IV-G5 Ratio of Important Tokens
@@ V Discussion
High-resolution image classification is not the only field that benefits from this memory efficiency. FPT+ can also be applied to ViT variants with smaller patch sizes. ViTs with smaller patch sizes generally have more powerful abilities to extract fine-grained semantics. However, smaller patch sizes result in longer input sequences for the same input resolution. FPT+ may be a good solution to attain the best balance. The memory efficiency of FPT+ also allows for larger batch sizes during training, which can accelerate the training process and potentially improve the stability and performance of transfer learning. Additionally, there are settings in natural image analysis that also require high resolution, such as satellite imagery [1]. The generalizability of FPT+ to natural images remains unexplored, which is one of our future works.
VI Conclusion
In this paper, we extend our previous work FPT [23] to FPT+. FPT+ significantly improves the memory efficiency over existing PETL methods, particularly in the high-resolution context commonly encountered in medical image analysis. To reduce memory consumption, we first enhance the design of side tuning with asymmetric input. To effectively adapt pre-trained knowledge from the LPM, we introduce fine-grained prompts and fusion modules to summarize and process intermediate features of the LPM and effectively convey them to the side network. To further reduce memory utilization, we propose the selection of important tokens and the preloading of pre-trained features. By integrating all these components, we present a parameter and memory-efficient transfer learning method for high-resolution medical image classification. The effectiveness of FPT+ is successfully established across eight medical image datasets of diverse sizes, modalities, and complexities. Experimental results show that FPT+ outperforms all other SOTA PETL methods while maintaining the best parameter and memory efficiency.